This problem says to complete the following table....

Line 1....[H3O+] [OH-] pH Acidic, Base or Neutral?
Line 2... ?? ?? 11.0 ??
Line 3... ?? ?? ?? Neutral
Line 4... ?? ?? 4.20 ??
Line 5. 7.0x10^-12M ?? 2.60 ??
1
Expert's answer
2015-11-19T11:35:42-0500
Line 5 – mistake. If you have pH 2.6, [H3O+] = 3.98 × 10^2 M and [OH-] = 2.51 × 10^11 M.
If [H3O+] = 7.0x10^-12 M, pH = 11.2 (Base)
pH = - lg [H3O+]